# Exploring sub-GeV dark matter via $s$-wave, $p$-wave, and resonance annihilation with CMB data

### v1

@@ Abstract
We revisit constraints on sub-GeV dark matter (DM) annihilation via $s$-wave, $p$-wave, and resonance processes using current and future CMB data from Planck, FIRAS, and upcoming experiments as LiteBird, CMB-S4, PRISTINE, and PIXIE. For $s$-wave annihilation, we provide updated limits for both $e^{+}e^{-}$ and $\pi\pi$ channels, with the profile likelihood method yielding stronger constraints than marginal posterior method. In the $p$-wave case, we comprehensively conduct a model-independent inequality for the 95% upper limits from FIRAS, PRISTINE, and PIXIE, with future experiments expected to surpass current BBN limits. For resonance annihilation, we report—for the first time—the $95\%$ upper limits on the decay branching ratio of the mediator particle, when resonances peak during the recombination epoch. Overall, our study highlights the complementary strengths of $\mu$-distortion and CMB anisotropies in probing sub-GeV DM annihilation.
@@ I Introduction
Dark matter (DM) is widely acknowledged as a fundamental constituent of the Universe and a primary agent in the formation of cosmological structures. However, its nature remains an elusive mystery. The observed DM relic density provides an important window into understanding the nature of DM. Among the various mechanisms proposed to account for the relic density, the weakly interacting massive particle as thermally produced via the freeze-out mechanism from the Standard Model (SM) plasma in the early Universe is the most compelling scenarios Goldberg (1983); Goodman and Witten (1985); Blumenthal et al. (1984) (see Jungman et al. (1996); Bergström (2000); Bertone et al. (2005) as reviews). Canonical models, such as supersymmetry, predict weakly interacting massive particles with masses around $\mathcal{O}(100)$ GeV Ellis et al. (1984), but decades of direct detection experiments, including XENON Aprile (2013); Aprile et al. (2023), PandaX Cui et al. (2017); Meng et al. (2021), DarkSide-50 Agnes et al. (2023), LZ Aalbers et al. (2023) and CDEX Jiang et al. (2018); Zhang et al. (2022), have yielded no conclusive signals, compressing the parameter space for massive DM models Akerib et al. (2022).
@@
The absence of discoveries has spurred critical reassessments of both experimental methodologies and theoretical assumptions on light DM consisting of particles in the MeV-GeV range. For light DM, an universal bound of light DM mass was originally studied by Lee and Weinberg Lee and Weinberg (1977) and concluded that if the mediator of the DM-SM interaction is parts of the SM particles, the DM mass would have to be greater than a lower bound of the order of $2\,\mathrm{GeV}$. To beat the bound, except for non-thermal freeze scenario Hall et al. (2010); Bernal et al. (2017); Dvorkin et al. (2021) and asymmetric DM models Kaplan et al. (2009); Zurek (2014), another typical way to escape the Lee-Weinberg constraints is to study light DM by introducing new “dark sector” particles, i.e., new auxiliary forces and matter fields other than DM. For a detailed review, we refer readers to Ref. Lin (2019). In seek of sub-GeV DM particles, indirect detection experiments leverage cosmic ray detectors DAMPE Chang et al. (2017), Fermi-LAT Ajello et al. (2021), and AMS Aguilar et al. (2013) and next generation gamma-ray telescopes including VLAST Pan et al. (2024), LASSHO Addazi et al. (2022), COSI Tomsick (2021), CTA Acharya et al. (2018), GECCO Moiseev (2023), and GRAMS Aramaki (2024) (see Bertuzzo et al. (2017); O’Donnell and Slatyer (2024) as reviews) to probe annihilation or decay signals in the Galactic Center and dwarf galaxies.
@@
The Cosmic Microwave Background (CMB) serves as a critical probe for MeV-GeV DM due to its sensitivity to energy injection during recombination and alterations in the relativistic energy density $N_{\rm eff}$ Sabti et al. (2020); Lin et al. (2012). In the thermal history of universe, DM annihilation generate high-energy photons and electrons that heat and ionize hydrogen and helium gas, broadening the last scattering surface. This results in a relative suppression to the temperature fluctuations and enhancement to the polarization Padmanabhan and Finkbeiner (2005a) of CMB. Meanwhile deviations of the CMB frequency spectrum named as blackbody spectral distortions of CMB Zeldovich and Sunyaev (1969); Sunyaev and Zeldovich (1970); Burigana et al. (1991); Hu and Silk (1993); Chluba and Sunyaev (2012); Chluba et al. (2012), which raise a complementary probes to study energy injection processes in the early universe. These limitations can be quantitatively expressed through bounds on the thermally-averaged $s$-wave annihilation cross-section $\langle\sigma v_{\text{rel}}\rangle$ into SM final states, with Planck satellite observations during the recombination epoch establishing an upper limit of $\langle\sigma v_{\text{rel}}\rangle_{\mathrm{CMB}}\lesssim 3\times 10^{-26}\mathrm{cm}^{3}\mathrm{s}^{-1}(m_{\chi}/10\,\mathrm{GeV})$ Aghanim et al. (2020a), where $m_{\chi}$ is DM mass. Given that standard thermal relic DM scenarios predict a thermally-averaged annihilation cross-section $\langle\sigma v_{\text{rel}}\rangle\sim 10^{-26}\mathrm{cm}^{3}\mathrm{s}^{-1}$ during the freeze-out epoch, pure $s$-wave annihilation scenario is excluded for DM masses $m_{\chi}\lesssim 10\,\mathrm{GeV}$. However, when accounting for velocity-dependent scattering cross-sections, the approximation $\langle\sigma v_{\text{rel}}\rangle\sim 10^{-26}\,\mathrm{cm}^{3}\,\mathrm{s}^{-1}$ no longer holds. Also, $\langle\sigma v_{\text{rel}}\rangle$ obtained from precise relic density calculations Chatterjee and Hryczuk (2025); Duan et al. (2024); Binder et al. (2017) cannot be directly related to $\langle\sigma v_{\text{rel}}\rangle_{\mathrm{CMB}}$. Furthermore, by incorporating velocity dependence, the annihilation rate during the recombination epoch can be suppressed due to the reduced relative velocities of DM particles, thereby evading constraints from CMB observations. Among these mechanisms, $p$-wave annihilation (velocity-squared dependent) Boehm and Fayet (2004); Kumar and Marfatia (2013) and resonant enhancement scenarios Ibe et al. (2009); Griest and Seckel (1991) have emerged as viable solutions in recent studies.
In this article, we investigate the constraints on sub-GeV DM annihilation imposed by the data of COBE/FAIRS Fixsen et al. (1996a); Mather et al. (1999), Planck satellite Aghanim et al. (2020a, b) and Baryon Acoustic Oscillations (BAO) Alam et al. (2017); Buen-Abad et al. (2018); Beutler et al. (2011); Ross et al. (2015), considering both $e^{+}e^{-}$ and $\pi\pi$ final states through $s$-wave, $p$-wave, and resonant processes.
@@
We utilize the Hazma tool Coogan et al. (2020a) to generate precise photon and electron spectra, and the energy deposition into the intergalactic medium (IGM) is calculated using DarkHistory v2.0 Liu et al. (2020). This systematic approach allows us to probe the nature of light DM and its interactions with greater precision, offering new insights into the phenomenology of sub-GeV DM annihilation.
@@ II.2 Energy Deposition
When DM particles annihilate into primary particles, such as SM particles and long-lived new particles, these can decay into stable particles and inject energy into the universe, while neutrino final states do not contribute to gas heating or ionization. For DM masses from a few MeVs to a few GeVs, possible annihilation channels include two-body final states (e.g., $2e^{\pm}$, $2\mu^{\pm}$, or $2\pi$) or four-body final states (e.g., $4e^{\pm}$, $4\mu^{\pm}$, or $4\pi$), which arise from the decay of two long-lived new particles. However, Planck constraints on four-body models are similar to those on two-body models across a wide DM mass range, as the constraints are mainly sensitive to the total energy of $e^{\pm}$ and $\gamma$ final states, not their spectrum shapes Clark et al. (2017). In this work, we focus on two-body final states (electron-positron pairs and pions), since electrons and positrons are leptons, while pions are mesons made of up and down quarks. For simplicity, we adopt the Higgs portal model to compute the pion branch ratios, $\text{BR}(\chi\chi\rightarrow\pi^{+}\pi^{-})=2/3$ and $\text{BR}(\chi\chi\rightarrow\pi^{0}\pi^{0})=1/3$.
We compute DM energy deposition energy that goes into ionization, heating and excitation of the IGM, as
@@ III.1 Blackbody spectral distortions
where $x\equiv E_{\gamma}/T_{z}$ and $T_{z}\equiv T_{\text{CMB,0}}(1+z)$ represent the reference temperature in equilibrium with the thermal bath [^1] Radiation actual temperature $T_{\gamma}$ may differ from $T_{z}$ when energy transfer happens during photon scattering with electrons. However, it is just the temperature deviation from the reference temperature, which is eliminated by coinciding the reference temperature at $z=0$ with the observation. In an expanding universe, above interactions weaken over time and become insufficient before recombination, allowing DM-induced energy injections to leave imprints on the CMB frequency spectrum Chluba (2013a); Lucca et al. (2020); Lucca (2023); Li (2024).
@@
Since the universe is fully ionized, and energy injection follows the on-the-spot approximation Padmanabhan and Finkbeiner (2005a). Under such an assumption, DM annihilation into photons and electrons results in instantaneous and complete energy deposition into heating the IGM.
@@
The FIRAS experiment was aboard the COBE satellite provided one of the most precise measurements of the CMB spectrum. It has reported $|\mu|<4.7\times 10^{-5}$ Bianchini and Fabbian (2022), improving the previous monopole $\mu$-distortion limit $|\mu|<9\times 10^{-5}$ Fixsen et al. (1996b), due to more robust foreground cleaning.
Future missions aim to refine these limits and potentially detect smaller distortions. The 95% upper limit for $\mu$ from PRISTINE will be $|\mu|<8\times 10^{-7}$ Chluba et al. (2019), while PIXIE (Primordial Inflation Explorer) will set a more stringent limit of $|\mu|<8\times 10^{-8}$ Chluba et al. (2019) at the same confidence level. [^2] Super-PIXIE allow for potential detection of $\mu$-distortion values as low as $|\mu|\approx 2\times 10^{-8}$ at $3\sigma$ confidence level Chluba et al. (2019); Chluba (2016), the constraint from Super-PIXIE is not considered in this studied.
@@ III.2 CMB anisotropies
spectrum because it enhances the visibility of reionization-induced polarization Padmanabhan and Finkbeiner (2005a).
DM annihilation produces high-energy photons and electrons that heat and ionize hydrogen gas, broadening the last scattering surface. This increases the residual free electron fraction after recombination, enhancing Thomson scattering. As a result, the damping tail of the CMB temperature power spectrum at small angular scales is suppressed due to photon redistribution Padmanabhan and Finkbeiner (2005b); Galli et al. (2009); Slatyer et al. (2009). Furthermore, increased ionization at later times boosts the low-$l$ polarization power spectrum by enhancing the visibility of reionization-induced polarization Padmanabhan and Finkbeiner (2005a).
@@ III.3 Statistical framework
In this work, we utilize two types of likelihood distributions in the analysis. For blackbody spectral distortions, we adopt the 95% upper limit, $|\mu|<4.7\times 10^{-5}$, from FIRAS, meaning that the likelihood probabilities for a predicted $|\mu|$ exceeding $4.7\times 10^{-5}$ are zero. Similarly, we apply upper limits to the likelihoods for all future experimental prospects, including the future measurements of blackbody spectral distortions and CMB anisotropies. For data measurements, such as the CMB angular power spectra (TT, TE, and EE), we use Gaussian likelihoods included in the numerical MontePython Brinckmann and Lesgourgues (2018). To complete the likelihoods for CMB anisotropies, we incorporate the Planck 2018 data Aghanim et al. (2020a), including $-\ell$ (TT, TE, EE) spectra, $-\ell$ temperature TT spectrum, $-\ell$ polarization EE spectrum, lensing power spectrum, and BAO data Alam et al. (2017); Buen-Abad et al. (2018); Beutler et al. (2011); Ross et al. (2015).
To present our results, we use two statistical methods: the “marginal posterior” (MP) method and the “profile likelihood” (PL) method. The MP approach integrates the probability density over nuisance parameters by marginalizing the posterior, commonly used in cosmology, especially within the $\Lambda$CDM framework. In contrast, the PL method is preferred for null signal searches due to large volume effects and prior dependencies in unconstrained likelihoods, and is widely applied in DM direct and indirect detection. For the MP method, we use the prior distributions as described in Ref. Lewis (2013) on six cosmological parameters as nuisances—$\{\Omega_{b}h^{2},\Omega_{\text{CDM}}h^{2},100\theta_{s},\log(10^{10}A_{s}),n_{s},\tau_{\text{reio}}\}$—informed by Planck precise central values and uncertainties. However, we apply log priors for $\langle\sigma v_{\text{rel}}\rangle$, $b$, and $B_{i}B_{f}$ for three scenarios, respectively. For the PL method, we perform several fine scans in addition to the original Bayesian scan to improve coverage. For both methods, we exclude parameter regions where the accumulated probability exceeds 95% for each fixed $m_{\chi}$.

### v2

@@ Abstract
We revisit constraints on sub-GeV dark matter (DM) annihilation via $s$-wave, $p$-wave, and resonance processes using current and future CMB data from Planck, FIRAS, and upcoming experiments such as LiteBIRD, CMB-S4, PRISTINE, and PIXIE. For $s$-wave annihilation, we provide updated limits for both $e^{+}e^{-}$ and $\pi\pi$ channels, with the profile likelihood method yielding stronger constraints than the marginal posterior method. In the $p$-wave case, we comprehensively present a model-independent inequality for the 95% upper limits from FIRAS, PRISTINE, and PIXIE, with future experiments expected to surpass current BBN limits. For resonance annihilation, we report—for the first time—the $95\%$ upper limits on the decay branching ratio of the mediator particle, when the resonance peaks during the recombination epoch. Overall, our study highlights the complementary strengths of $\mu$-distortion and CMB anisotropies in probing sub-GeV DM annihilation.
@@ I Introduction
Dark matter (DM) is widely acknowledged as a fundamental constituent of the Universe and a primary agent in the formation of cosmological structures. However, its nature remains an elusive mystery. The observed DM relic density provides an important window into understanding the nature of DM. Among the various mechanisms proposed to account for the relic density, the weakly interacting massive particle as thermally produced via the freeze-out mechanism from the Standard Model (SM) plasma in the early Universe is the most compelling scenario Goldberg (1983); Goodman and Witten (1985); Blumenthal et al. (1984) (see Jungman et al. (1996); Bergström (2000); Bertone et al. (2005) as reviews). Canonical models, such as supersymmetry, predict weakly interacting massive particles with masses around $\mathcal{O}(100)$ GeV Ellis et al. (1984), but decades of direct detection experiments, including XENON Aprile (2013); Aprile et al. (2023), PandaX Cui et al. (2017); Meng et al. (2021), DarkSide-50 Agnes et al. (2023), LZ Aalbers et al. (2023) and CDEX Jiang et al. (2018); Zhang et al. (2022), have yielded no conclusive signals, compressing the parameter space for massive DM models Akerib et al. (2022).
@@
The absence of discoveries has spurred critical reassessments of both experimental methodologies and theoretical assumptions on light DM consisting of particles in the MeV-GeV range. For light DM, a universal bound of light DM mass was originally studied by Lee and Weinberg Lee and Weinberg (1977) and concluded that if the mediator of the DM-SM interaction is part of the SM particles, the DM mass would have to be greater than a lower bound of the order of $2\,\mathrm{GeV}$. To beat the bound, except for non-thermal freeze-out mechanisms Hall et al. (2010); Bernal et al. (2017); Dvorkin et al. (2021) and asymmetric DM models Kaplan et al. (2009); Zurek (2014), another typical way to escape the Lee-Weinberg constraints is to study light DM by introducing new “dark sector” particles, i.e., new force mediators and matter fields other than DM. For a detailed review, we refer readers to Ref. Lin (2019). To search for sub-GeV DM particles with masses above the MeV scale, indirect detection experiments utilize the cosmic ray detector Voyager-1 Stone et al. (2013) and various gamma-ray telescopes, including Fermi-LAT Ajello et al. (2021), COMPTEL Strong et al. (1994), INTEGRAL-SPI Siegert et al. (2015), VLAST Pan et al. (2024), GECCO Moiseev (2023), and GRAMS Aramaki (2024) (see Bertuzzo et al. (2017); O’Donnell and Slatyer (2024) for reviews) to probe annihilation or decay signals in the Galactic Center and dwarf galaxies De la Torre Luque et al. (2024a, b); Cirelli et al. (2021).
@@
The Cosmic Microwave Background (CMB) serves as a critical probe for MeV-GeV DM due to its sensitivity to energy injection during recombination Galli et al. (2013); Slatyer and Wu (2017); Zhang et al. (2007); Chen and Kamionkowski (2004); Galli et al. (2009); Slatyer et al. (2009); Poulin et al. (2017); Padmanabhan and Finkbeiner (2005); Kanzaki et al. (2010); Galli et al. (2011); Diamanti et al. (2013) and alterations in the relativistic energy density $N_{\rm eff}$ Sabti et al. (2020); Lin et al. (2012). In the thermal history of the universe, DM annihilation generates high-energy photons and electrons that heat and ionize hydrogen and helium gas, broadening the last scattering surface. This results in a relative suppression of temperature fluctuations accompanied by an enhancement of polarization in the CMB Padmanabhan and Finkbeiner (2005). Meanwhile, deviations of the CMB frequency spectrum, known as blackbody spectral distortions of CMB Zeldovich and Sunyaev (1969); Sunyaev and Zeldovich (1970); Burigana et al. (1991); Hu and Silk (1993); Chluba and Sunyaev (2012); Chluba et al. (2012), provide a complementary probe to study energy injection processes in the early universe. These effects can be quantitatively constrained through bounds on the thermally-averaged $s$-wave annihilation cross-section $\langle\sigma v_{\text{rel}}\rangle$ into SM final states. Observations from the Planck satellite during the recombination epoch establish a limit of $\langle\sigma v_{\text{rel}}\rangle_{\mathrm{CMB}}\lesssim 3\times 10^{-26}\mathrm{cm}^{3}\mathrm{s}^{-1}(m_{\chi}/10\,\mathrm{GeV})$ Aghanim et al. (2020a), where $m_{\chi}$ is DM mass. Given that standard thermal relic DM scenarios predict a thermally-averaged annihilation cross-section $\langle\sigma v_{\text{rel}}\rangle\sim 10^{-26}\mathrm{cm}^{3}\mathrm{s}^{-1}$ during the freeze-out epoch, pure $s$-wave annihilation scenario is excluded for DM masses $m_{\chi}\lesssim 10\,\mathrm{GeV}$. However, when accounting for velocity-dependent annihilation cross-sections, the approximation $\langle\sigma v_{\text{rel}}\rangle\sim 10^{-26}\,\mathrm{cm}^{3}\,\mathrm{s}^{-1}$ no longer holds. Also, $\langle\sigma v_{\text{rel}}\rangle$ obtained from precise relic density calculations Chatterjee and Hryczuk (2025); Duan et al. (2024); Binder et al. (2017) cannot be directly related to $\langle\sigma v_{\text{rel}}\rangle_{\mathrm{CMB}}$. Furthermore, by incorporating velocity dependence, the annihilation rate during the recombination epoch can be suppressed due to the reduced relative velocities of DM particles, thereby evading constraints from CMB observations. Among these mechanisms, $p$-wave annihilation (velocity-squared dependent) An et al. (2017); Boehm and Fayet (2004); Kumar and Marfatia (2013) and resonant enhancement scenarios Ibe et al. (2009); Griest and Seckel (1991) have emerged as viable solutions in recent studies.
@@
In this article, we investigate the constraints on sub-GeV DM annihilation imposed by the data of COBE/FIRAS Fixsen et al. (1996a); Mather et al. (1999), Planck satellite Aghanim et al. (2020a, b) and Baryon Acoustic Oscillations (BAO) Alam et al. (2017); Buen-Abad et al. (2018); Beutler et al. (2011); Ross et al. (2015), considering both $e^{+}e^{-}$ and $\pi\pi$ final states through $s$-wave, $p$-wave, and resonant processes. We utilize the Hazma tool Coogan et al. (2020a) to generate precise photon and electron spectra, and calculate the energy deposition into the intergalactic medium (IGM) using DarkHistory v2.0 Liu et al. (2020). This systematic approach allows us to probe the nature of light DM and its interactions with greater precision, offering new insights into the phenomenology of sub-GeV DM annihilation.
@@ II.2 Energy Deposition
When DM particles annihilate into primary particles, such as SM particles and long-lived new particles, these can decay into stable particles and inject energy into the universe, while neutrino final states do not contribute to gas heating or ionization. For DM masses from a few MeV to a few GeV, possible annihilation channels include two-body final states (e.g., $2e^{\pm}$, $2\mu^{\pm}$, or $2\pi$) or four-body final states (e.g., $4e^{\pm}$, $4\mu^{\pm}$, or $4\pi$), which arise from the decay of two long-lived new particles. However, Planck constraints on four-body models are similar to those on two-body models across a wide DM mass range, as the constraints are mainly sensitive to the total energy of $e^{\pm}$ and $\gamma$ final states, not their spectrum shapes Clark et al. (2017). In this work, we focus on two-body final states (electron-positron pairs and pions), since electrons and positrons are leptons, while pions are mesons made of up and down quarks. For simplicity, we adopt the Higgs portal model to compute the pion branch ratios, $\text{BR}(\chi\chi\rightarrow\pi^{+}\pi^{-})=2/3$ and $\text{BR}(\chi\chi\rightarrow\pi^{0}\pi^{0})=1/3$.
We compute DM energy deposition that goes into ionization, heating and excitation of the IGM, as
@@ III.1 Blackbody spectral distortions
where $x\equiv E_{\gamma}/T_{z}$ and $T_{z}\equiv T_{\text{CMB,0}}(1+z)$ represent the reference temperature in equilibrium with the thermal bath. [^1] Radiation actual temperature $T_{\gamma}$ may differ from $T_{z}$ when energy transfer happens during photon scattering with electrons. However, it is just the temperature deviation from the reference temperature, which is eliminated by coinciding the reference temperature at $z=0$ with the observation. In an expanding universe, above interactions weaken over time and become insufficient before recombination, allowing DM-induced energy injections to leave imprints on the CMB frequency spectrum Chluba (2013a); Lucca et al. (2020); Lucca (2023); Li (2024).
@@
Since the universe is fully ionized, and the energy injection follows the on-the-spot approximation Padmanabhan and Finkbeiner (2005); Slatyer et al. (2009). Under such an assumption, DM annihilation into photons and electrons results in instantaneous and complete energy deposition into heating the IGM.
@@
The FIRAS experiment aboard the COBE satellite provided one of the most precise measurements of the CMB spectrum. It reported $|\mu|<4.7\times 10^{-5}$ Bianchini and Fabbian (2022), improving the previous monopole $\mu$-distortion limit $|\mu|<9\times 10^{-5}$ Fixsen et al. (1996b), due to more robust foreground cleaning.
Future missions aim to refine these limits and potentially detect smaller distortions. The 95% upper limit for $\mu$ from PRISTINE will be $|\mu|<8\times 10^{-7}$ Chluba et al. (2019), while PIXIE will set a more stringent limit of $|\mu|<8\times 10^{-8}$ Chluba et al. (2019) at the same confidence level. [^2] Super-PIXIE allows for potential detection of $\mu$-distortion values as low as $|\mu|\approx 2\times 10^{-8}$ at $3\sigma$ confidence level Chluba et al. (2019); Chluba (2016); the constraint from Super-PIXIE is not considered in this study.
@@ III.2 CMB anisotropies
DM annihilation produces high-energy photons and electrons that heat and ionize hydrogen gas, broadening the last scattering surface. This increases the residual free electron fraction after recombination, enhancing Thomson scattering. As a result, the damping tail of the CMB temperature power spectrum at small angular scales is suppressed due to photon redistribution Padmanabhan and Finkbeiner (2005); Galli et al. (2009); Slatyer et al. (2009). Furthermore, increased ionization at later times boosts the low-$l$ polarization power spectrum by enhancing the visibility of reionization-induced polarization Padmanabhan and Finkbeiner (2005).
@@ III.3 Statistical framework
In this work, we utilize two types of likelihood functions in the analysis. For blackbody spectral distortions, we adopt the 95% upper limit, $|\mu|<4.7\times 10^{-5}$, from FIRAS, meaning that the likelihood probabilities for a predicted $|\mu|$ exceeding $4.7\times 10^{-5}$ are zero. Similarly, we apply upper limits to the likelihoods for all future experimental prospects, including the future measurements of blackbody spectral distortions and CMB anisotropies. For data measurements, such as the CMB angular power spectra (TT, TE, and EE), we use Gaussian likelihoods included in the numerical MontePython Brinckmann and Lesgourgues (2018). To complete the likelihoods for CMB anisotropies, we incorporate the Planck 2018 data Aghanim et al. (2020a), including $-\ell$ (TT, TE, EE) spectra, $-\ell$ temperature TT spectrum, $-\ell$ polarization EE spectrum, lensing power spectrum, and BAO data Alam et al. (2017); Buen-Abad et al. (2018); Beutler et al. (2011); Ross et al. (2015).
To present our results, we use two statistical methods: the “marginal posterior” (MP) method and the “profile likelihood” (PL) method. The MP approach integrates the probability density over nuisance parameters by marginalizing the posterior, commonly used in cosmology, especially within the $\Lambda$CDM framework. In contrast, the PL method is preferred for null signal searches due to parameter volume effects and prior dependencies in unconstrained likelihoods, and is widely applied in DM direct and indirect detection. For the MP method, we use the prior distributions as described in Ref. Lewis (2013) on six cosmological parameters as nuisances—$\{\Omega_{b}h^{2},\Omega_{\text{CDM}}h^{2},100\theta_{s},\log(10^{10}A_{s}),n_{s},\tau_{\text{reio}}\}$—informed by Planck precise central values and uncertainties. However, we assign log-uniform priors for $\langle\sigma v_{\text{rel}}\rangle$, $b$, and $B_{i}B_{f}$ for three scenarios, respectively. For the PL method, we perform several fine scans in addition to the original Bayesian scan to improve coverage. For both methods, we exclude parameter regions where the accumulated probability exceeds 95% for each fixed $m_{\chi}$.